# Semi-Supervised Deep Sobolev Regression: Estimation, Variable Selection and Beyond

### v1

@@ D.2 Semi-Supervised Deep Sobolev Regressor
$\leq B_{1,k}^{2}\Big{(}\frac{2\log N(\delta,\operatorname{conv}(D_{k}{\mathcal{F}}),L^{2}({\mathcal{S}}))}{m}\Big{)}^{1/2}+B_{1,k}\delta.$
By using Lemma A.5, we obtain that for $1\leq k\leq d$,
$\leq c\Big{\{}B_{1,k}^{2}\Big{(}\frac{\operatorname{VCdim}(D_{k}{\mathcal{F}})}{m\delta}\Big{)}^{1/2}+B_{1,k}\delta\Big{\}},$
where $c$ is an absolute positive constant. Summing over this equation for $1\leq k\leq d$ yields
${\mathbb{E}}_{{\mathcal{S}}}\Big{[}(\nabla f_{0},\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0}))_{L^{2}(\nu_{X})}-(\nabla f_{0},\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0}))_{L^{2}({\mathcal{S}})}\Big{]}$
$\leq c\inf_{\delta>0}\Big{\{}B_{1}^{2}\Big{(}\frac{\operatorname{VCdim}(\nabla{\mathcal{F}})}{m\delta}\Big{)}^{1/2}+B_{1}\delta\Big{\}}.$
Combining this with Lemma A.1 and Assumption 4, we find that
@@
$\leq\lambda{\mathbb{E}}_{{\mathcal{S}}}\Big{[}-(\nabla f_{0},\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f))_{L^{2}(\nu_{X})}\Big{]}+c\lambda\inf_{\delta>0}\Big{\{}B_{1}^{2}\Big{(}\frac{\operatorname{VCdim}(\nabla{\mathcal{F}})}{m\delta}\Big{)}^{1/2}+B_{1}\delta\Big{\}}$
$=\lambda{\mathbb{E}}_{{\mathcal{S}}}\Big{[}(\Delta f_{0},\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f)_{L^{2}(\nu_{X})}+(\nabla f_{0}\cdot\nabla(\log q),\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f)_{L^{2}(\nu_{X})}\Big{]}$
$\quad+c\lambda\inf_{\delta>0}\Big{\{}B_{1}^{2}\Big{(}\frac{\operatorname{VCdim}(\nabla{\mathcal{F}})}{m\delta}\Big{)}^{1/2}+B_{1}\delta\Big{\}}$
$\leq\lambda\kappa^{1/2}{\mathbb{E}}_{{\mathcal{S}}}\Big{[}\Big{\{}\|\Delta f_{0}\|_{L^{2}(\nu_{X})}+\|\nabla f_{0}\cdot\nabla(\log q)\|_{L^{2}(\nu_{X})}\Big{\}}\Big{\{}R(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})^{1/2}+\|f-f_{0}\|_{L^{2}(\mu_{X})}\Big{\}}\Big{]}$
$\quad+c\lambda\inf_{\delta>0}\Big{\{}B_{1}^{2}\Big{(}\frac{\operatorname{VCdim}(\nabla{\mathcal{F}})}{m\delta}\Big{)}^{1/2}+B_{1}\delta\Big{\}}$
$\leq 9\lambda^{2}\kappa\Big{\{}\|\Delta f_{0}\|_{L^{2}(\nu_{X})}^{2}+\|\nabla f_{0}\cdot\nabla(\log q)\|_{L^{2}(\nu_{X})}^{2}\Big{\}}+\frac{1}{16}{\mathbb{E}}_{{\mathcal{S}}}\Big{[}R(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})\Big{]}+\frac{1}{2}\|f-f_{0}\|_{L^{2}(\mu_{X})}^{2}$
$\quad+\lambda^{2}B_{1}^{2}+c^{2}\frac{B_{1}^{2}}{2}\frac{\operatorname{VCdim}(\nabla{\mathcal{F}})}{m\delta}+\delta^{2},$
@@
$=\lambda(\nabla\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda},\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f))_{L^{2}({\mathcal{S}})}+\frac{1}{n}\sum_{i=1}^{n}(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}(X_{i})-Y_{i})(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}(X_{i})-f(X_{i}))\leq 0.$
@@
$\lambda(\nabla\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda},\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f))_{L^{2}({\mathcal{S}})}+(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0},\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f)_{L^{2}({\mathcal{D}})}$ (55)
@@
$=\lambda\|\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0})\|_{L^{2}({\mathcal{S}})}^{2}-\lambda(\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0}),\nabla(f-f_{0}))_{L^{2}({\mathcal{S}})}$
$\quad+\lambda(\nabla f_{0},\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f))_{L^{2}({\mathcal{S}})},$ (57)
@@
$\lambda(\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0}),\nabla(f-f_{0}))_{L^{2}({\mathcal{S}})}\leq\frac{\lambda}{2}\|\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0})\|_{L^{2}({\mathcal{S}})}^{2}+\frac{\lambda}{2}\|\nabla(f-f_{0})\|_{L^{2}({\mathcal{S}})}^{2}.$ (58)
@@
$\leq\lambda{\mathbb{E}}_{{\mathcal{S}}}\Big{[}(\nabla\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda},\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f))_{L^{2}({\mathcal{S}})}\Big{]}+\frac{\lambda}{2}\|\nabla(f-f_{0})\|_{L^{2}(\nu_{X})}^{2}$
@@
$\leq\lambda{\mathbb{E}}_{{\mathcal{S}}}\Big{[}(\nabla\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda},\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f))_{L^{2}({\mathcal{S}})}\Big{]}+\frac{\lambda}{2}\|\nabla(f-f_{0})\|_{L^{2}(\nu_{X})}^{2}$
$\quad+\frac{1}{16}{\mathbb{E}}_{{\mathcal{S}}}\Big{[}R(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})\Big{]}+c\Big{\{}\tilde{\beta}\lambda^{2}+\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}({\mathcal{F}},m)\Big{\}}.$
@@
$(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0},\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f)_{L^{2}({\mathcal{D}})}$ $=\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})-(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0},f-f_{0})_{L^{2}({\mathcal{D}})}$
@@
$\frac{1}{2}{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})\Big{]}\leq\frac{1}{2}R(f)+{\mathbb{E}}_{{\mathcal{D}}}\Big{[}(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0},\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f)_{L^{2}({\mathcal{D}})}\Big{]}.$ (60)
@@
${\mathbb{E}}_{{\mathcal{D}},{\mathcal{S}}}\Big{[}\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})\Big{]}$
$\leq\frac{1}{4}{\mathbb{E}}_{{\mathcal{D}},{\mathcal{S}}}\Big{[}R(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})\Big{]}+c\Big{\{}\tilde{\beta}\lambda^{2}+\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)+\varepsilon_{\operatorname{gen}}({\mathcal{F}},n)+\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}({\mathcal{F}},m)\Big{\}}.$
Then according to (29), it follows that
${\mathbb{E}}_{{\mathcal{D}},{\mathcal{S}}}\Big{[}R(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})\Big{]}\leq 2{\mathbb{E}}_{{\mathcal{D}},{\mathcal{S}}}\Big{[}\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})\Big{]}+c^{\prime}B_{0}^{2}\frac{\operatorname{VCdim}({\mathcal{F}})}{n\log^{-1}(n)}$
$\leq\frac{1}{2}{\mathbb{E}}_{{\mathcal{D}},{\mathcal{S}}}\Big{[}R(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})\Big{]}+2c\Big{\{}\tilde{\beta}\lambda^{2}+\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)+\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}({\mathcal{F}},m)\Big{\}}+(2c+c^{\prime})\varepsilon_{\operatorname{gen}}({\mathcal{F}},n),$
@@
${\mathbb{E}}_{{\mathcal{D}},{\mathcal{S}}}\Big{[}\|\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0}\|_{L^{2}(\mu_{X})}^{2}\Big{]}\lesssim\tilde{\beta}\lambda^{2}+\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)+\varepsilon_{\operatorname{gen}}({\mathcal{F}},n)+\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}({\mathcal{F}},m).$ (61)
@@
Proof of Theorem 5.6.
$\|f-\phi\|_{L^{2}(\mu_{X})}\leq CN^{-s}\quad\text{and}\quad\|\nabla(f-\phi)\|_{L^{2}(\nu_{X})}\leq CN^{-(s-1)}.$
By using Lemma A.6, we find $\operatorname{VCdim}({\mathcal{F}})\leq cN^{d}\log N$. Further, applying Lemma A.7 deduces $\operatorname{VCdim}(D_{k}{\mathcal{F}})\leq cL^{2}S\leq N^{d}(\log N)^{2}$ for each $1\leq k\leq d$. Substituting these estimates into Lemma 5.5 yields
@@
$\lesssim\tilde{\beta}\lambda^{2}+CN^{-2s}+C\lambda N^{-2(s-1)}+C\inf_{\delta_{0}>0}\Big{\{}(B_{0}^{2}+\sigma^{2})\frac{N^{d}\log N}{n\delta_{0}}+(B_{0}+\sigma)\delta_{0}\Big{\}}$
$\quad+C\inf_{\delta_{1}>0}\Big{\{}B_{1}^{2}\frac{N^{d}(\log N)^{3}}{m\delta_{1}}+B_{1}\delta_{1}\Big{\}}.$
Setting $N=(1+\sigma)^{-\frac{3}{d+4s}}n^{\frac{1}{d+4s}}$, $\delta_{0}=\sigma^{\frac{1}{2}}N^{\frac{d}{2}}n^{-\frac{1}{2}}$, $\delta_{1}=\sigma^{\frac{1}{2}}N^{\frac{d}{2}}m^{-\frac{1}{2}}$ and letting the regularization parameter be $\lambda=(1+\sigma)^{\frac{3s}{d+4s}}n^{-\frac{s}{d+4s}}$deduce the desired result. ∎
Appendix E Proofs in Results in Section 6
Proof of Corollary 6.3.
By Markov’s inequality (Mohri, Rostamizadeh and Talwalkar, 2018, Theorem C.11), the following inequality holds for each $\epsilon>0$
$\lim_{n\rightarrow\infty}\operatorname{Pr}\Big{\{}\|D_{k}\widehat{f}_{{\mathcal{D}}}^{\lambda}-D_{k}f_{0}\|_{L^{2}(\nu_{X})}>\epsilon\Big{\}}\leq\lim_{n\rightarrow\infty}\frac{{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\|D_{k}\widehat{f}_{{\mathcal{D}}}^{\lambda}-D_{k}f_{0}\|_{L^{2}(\nu_{X})}\Big{]}}{\epsilon}=0,$ (62)
where the equality follows from Corollary 6.2. For each irrelevant variable $k\notin{\mathcal{I}}(f_{0})$, one has $\|D_{k}f_{0}\|_{L^{2}(\nu_{X})}=0$. Then (62) deduces that for each $\epsilon>0$
$\lim_{n\rightarrow\infty}\operatorname{Pr}\Big{\{}\|D_{k}\widehat{f}_{{\mathcal{D}}}^{\lambda}\|_{L^{2}(\nu_{X})}>\epsilon\Big{\}}\leq\lim_{n\rightarrow\infty}\operatorname{Pr}\Big{\{}\|D_{k}\widehat{f}_{{\mathcal{D}}}^{\lambda}-D_{k}f_{0}\|_{L^{2}(\nu_{X})}>\epsilon\Big{\}}=0,$

### v2

@@ D-B Semi-Supervised Deep Sobolev Regressor
$\leq B_{1,k}^{2}\Big{(}\frac{2\log N(B_{1,k}\delta,D_{k}{\mathcal{F}},L^{2}({\mathcal{S}}))}{m}\Big{)}^{1/2}+B_{1,k}^{2}\delta.$
Summing over this equation for $1\leq k\leq d$ yields
${\mathbb{E}}_{{\mathcal{S}}}\Big{[}(\nabla f_{0},\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0}))_{L^{2}(\nu_{X})}$
$\quad-(\nabla f_{0},\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0}))_{L^{2}({\mathcal{S}})}\Big{]}$
$\leq\sum_{k=1}^{d}B_{1,k}^{2}\inf_{\delta>0}\Big{\{}\Big{(}\frac{2\log N(B_{1,k}\delta,D_{k}{\mathcal{F}},L^{2}({\mathcal{S}}))}{m}\Big{)}^{1/2}+\delta\Big{\}}.$
Combining this with Lemma A.1 and Assumption 4, we find that for each $\delta>0$,
@@
$\leq\lambda{\mathbb{E}}_{{\mathcal{S}}}\Big{[}-(\nabla f_{0},\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f))_{L^{2}(\nu_{X})}\Big{]}$
$\quad+\lambda\sum_{k=1}^{d}B_{1,k}^{2}\Big{\{}\Big{(}\frac{2\log N(B_{1,k}\delta,D_{k}{\mathcal{F}},L^{2}({\mathcal{S}}))}{m}\Big{)}^{1/2}+\delta\Big{\}}$
$=\lambda{\mathbb{E}}_{{\mathcal{S}}}\Big{[}(\Delta f_{0},\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f)_{L^{2}(\nu_{X})}$
$\quad+(\nabla f_{0}\cdot\nabla(\log q),\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f)_{L^{2}(\nu_{X})}\Big{]}$
$\quad+\lambda\sum_{k=1}^{d}B_{1,k}^{2}\Big{\{}\Big{(}\frac{2\log N(B_{1,k}\delta,D_{k}{\mathcal{F}},L^{2}({\mathcal{S}}))}{m}\Big{)}^{1/2}+\delta\Big{\}}$
$\leq\lambda\kappa^{1/2}{\mathbb{E}}_{{\mathcal{S}}}\Big{[}\Big{\{}\|\Delta f_{0}+\nabla f_{0}\cdot\nabla(\log q)\|_{L^{2}(\nu_{X})}\Big{\}}$
×$\quad\times\Big{\{}R(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})^{1/2}+\|f-f_{0}\|_{L^{2}(\mu_{X})}\Big{\}}\Big{]}$
$\quad+\lambda\sum_{k=1}^{d}B_{1,k}^{2}\Big{\{}\Big{(}\frac{2\log N(B_{1,k}\delta,D_{k}{\mathcal{F}},L^{2}({\mathcal{S}}))}{m}\Big{)}^{1/2}+\delta\Big{\}}$
$\leq 9\lambda^{2}\kappa\Big{\{}\|\Delta f_{0}\|_{L^{2}(\nu_{X})}^{2}+\|\nabla f_{0}\cdot\nabla(\log q)\|_{L^{2}(\nu_{X})}^{2}\Big{\}}$
$\quad+\frac{1}{16}{\mathbb{E}}_{{\mathcal{S}}}\Big{[}R(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})\Big{]}+\frac{1}{2}\|f-f_{0}\|_{L^{2}(\mu_{X})}^{2}+\lambda^{2}\Big{(}\sum_{k=1}^{d}B_{1,k}^{2}\Big{)}$
$\quad+\frac{1}{4}\sum_{k=1}^{d}B_{1,k}^{2}\Big{\{}\max_{1\leq k\leq d}\frac{2\log N(B_{1,k}\delta,D_{k}{\mathcal{F}},L^{2}({\mathcal{S}}))}{m}+\delta^{2}\Big{\}},$
@@
$=\lambda(\nabla\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda},\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f))_{L^{2}({\mathcal{S}})}$
$\quad+\frac{1}{n}\sum_{i=1}^{n}(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}(X_{i})-Y_{i})(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}(X_{i})-f(X_{i}))\leq 0.$
@@
$\quad+(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0},\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f)_{L^{2}({\mathcal{D}})}$
@@
$\lambda(\nabla\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda},\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f))_{L^{2}({\mathcal{S}})}$
@@
$=\lambda\|\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0})\|_{L^{2}({\mathcal{S}})}^{2}+\lambda(\nabla f_{0},\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f))_{L^{2}({\mathcal{S}})}$
$\quad-\lambda(\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0}),\nabla(f-f_{0}))_{L^{2}({\mathcal{S}})},$ (62)
@@
$\lambda(\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0}),\nabla(f-f_{0}))_{L^{2}({\mathcal{S}})}\\ \leq\frac{\lambda}{2}\|\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0})\|_{L^{2}({\mathcal{S}})}^{2}+\frac{\lambda}{2}\|\nabla(f-f_{0})\|_{L^{2}({\mathcal{S}})}^{2}.$ (63)
@@
$\leq\lambda{\mathbb{E}}_{{\mathcal{S}}}\Big{[}(\nabla\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda},\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f))_{L^{2}({\mathcal{S}})}\Big{]}$
$\quad+\frac{\lambda}{2}\|\nabla(f-f_{0})\|_{L^{2}(\nu_{X})}^{2}$
@@
$\leq\lambda{\mathbb{E}}_{{\mathcal{S}}}\Big{[}(\nabla\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda},\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f))_{L^{2}({\mathcal{S}})}\Big{]}$
$\quad+\frac{\lambda}{2}\|\nabla(f-f_{0})\|_{L^{2}(\nu_{X})}^{2}+\frac{1}{16}{\mathbb{E}}_{{\mathcal{S}}}\Big{[}R(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})\Big{]}$
$\quad+c\Big{\{}\tilde{\beta}\lambda^{2}+\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}(\nabla{\mathcal{F}},m)\Big{\}}.$
@@
$(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0},\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f)_{L^{2}({\mathcal{D}})}$
$=\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})-(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0},f-f_{0})_{L^{2}({\mathcal{D}})}$
@@
$\frac{1}{2}{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})\Big{]}\\ \leq\frac{1}{2}R(f)+{\mathbb{E}}_{{\mathcal{D}}}\Big{[}(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0},\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f)_{L^{2}({\mathcal{D}})}\Big{]}.$ (65)
@@
${\mathbb{E}}_{{\mathcal{D}},{\mathcal{S}}}\Big{[}\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})\Big{]}\leq\frac{1}{4}{\mathbb{E}}_{{\mathcal{D}},{\mathcal{S}}}\Big{[}R(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})\Big{]}\\ +c\Big{\{}\tilde{\beta}\lambda^{2}+\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)+\varepsilon_{\operatorname{gen}}({\mathcal{F}},n)+\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}(\nabla{\mathcal{F}},m)\Big{\}}.$
Then according to the above inequality and (39), it follows that
${\mathbb{E}}_{{\mathcal{D}},{\mathcal{S}}}\Big{[}R(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})\Big{]}\leq{\mathbb{E}}_{{\mathcal{D}},{\mathcal{S}}}\Big{[}\widehat{R}_{{\mathcal{D}}}(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})\Big{]}+c^{\prime}\varepsilon_{\operatorname{gen}}({\mathcal{F}},n)$
$\leq\frac{1}{2}{\mathbb{E}}_{{\mathcal{D}},{\mathcal{S}}}\Big{[}R(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda})\Big{]}+(2c+c^{\prime})\varepsilon_{\operatorname{gen}}({\mathcal{F}},n)$
$\quad+2c\Big{\{}\tilde{\beta}\lambda^{2}+\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)+\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}(\nabla{\mathcal{F}},m)\Big{\}},$
@@
${\mathbb{E}}_{{\mathcal{D}},{\mathcal{S}}}\Big{[}\|\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0}\|_{L^{2}(\mu_{X})}^{2}\Big{]}\\ \lesssim\tilde{\beta}\lambda^{2}+\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)+\varepsilon_{\operatorname{gen}}({\mathcal{F}},n)+\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}(\nabla{\mathcal{F}},m).$ (66)
@@
Proof of Theorem V.6.
According to Lemma V.3, we set the hypothesis class ${\mathcal{F}}$ as ReQU neural networks ${\mathcal{F}}={\mathcal{N}}(L,S)$ with $L={\mathcal{O}}(\log N)$ and $S={\mathcal{O}}(N^{d})$. Then there exists $f\in{\mathcal{F}}$ such that
$\|f-\phi\|_{L^{2}(\mu_{X})}\leq CN^{-s},\quad\|\nabla(f-\phi)\|_{L^{2}(\nu_{X})}\leq CN^{-(s-1)}.$
By using Lemma A.7 and set $\delta=1/n$, we find
$\log N(B_{1,k}n^{-1},D_{k}{\mathcal{F}},L^{2}({\mathcal{D}}))\\ \lesssim L^{2}S\log S\log n\lesssim N^{d}\log^{3}N\log n.$
Substituting these estimates and (54) into Lemma V.2 yields
@@
$\lesssim\tilde{\beta}\lambda^{2}+CN^{-2s}+C\lambda N^{-2(s-1)}$
$\quad+C\log n\Big{(}\frac{N^{d}\log N\log n}{n}\Big{)}^{\frac{1}{2}}+C\log n\frac{N^{d}\log^{3}N}{m}.$
Setting $N={\mathcal{O}}(n^{\frac{1}{d+4s}})$, and letting the regularization parameter be $\lambda={\mathcal{O}}(n^{-\frac{s}{d+4s}}\log^{2}n)$ deduce the desired result. ∎
Appendix E Proofs in Results in Section VI
Proof of Corollary VI.1.
A direct conclusion of Theorem V.4. ∎
Proof of Corollary VI.2.
By Markov’s inequality [97, Theorem C.11], the following inequality holds for each $\epsilon>0$
$\lim_{n\rightarrow\infty}\operatorname{Pr}\Big{\{}\|D_{k}\widehat{f}_{{\mathcal{D}}}^{\lambda}-D_{k}f_{0}\|_{L^{2}(\nu_{X})}>\epsilon\Big{\}}$ (67)
$\leq\lim_{n\rightarrow\infty}\frac{{\mathbb{E}}_{{\mathcal{D}}}\Big{[}\|D_{k}\widehat{f}_{{\mathcal{D}}}^{\lambda}-D_{k}f_{0}\|_{L^{2}(\nu_{X})}\Big{]}}{\epsilon}=0,$
where the equality follows from Corollary VI.1. For each irrelevant variable $k\notin{\mathcal{I}}(f_{0})$, one has $\|D_{k}f_{0}\|_{L^{2}(\nu_{X})}=0$. Then (67) deduces that for each $\epsilon>0$
$\lim_{n\rightarrow\infty}\operatorname{Pr}\Big{\{}\|D_{k}\widehat{f}_{{\mathcal{D}}}^{\lambda}\|_{L^{2}(\nu_{X})}>\epsilon\Big{\}}$